# Reinforcing Trustworthiness in Multimodal Emotional Support Systems

### v1

@@ Abstract
In today’s world, emotional support is increasingly essential, yet it remains challenging for both those seeking help and those offering it. Multimodal approaches to emotional support show great promise by integrating diverse data sources to provide empathetic, contextually relevant responses, fostering more effective interactions. However, current methods have notable limitations, often relying solely on text or converting other data types into text, or providing emotion recognition only, thus overlooking the full potential of multimodal inputs. Moreover, many studies prioritize response generation without accurately identifying critical emotional support elements or ensuring the reliability of outputs. To overcome these issues, we introduce MultiMood, a new framework that (i) leverages multimodal embeddings from video, audio, and text to predict emotional components and to produce responses responses aligned with professional therapeutic standards. To improve trustworthiness, we (ii) incorporate novel psychological criteria and apply Reinforcement Learning (RL) to optimize large language models (LLMs) for consistent adherence to these standards. We also (iii) analyze several advanced LLMs to assess their multimodal emotional support capabilities. Experimental results show that MultiMood achieves state-of-the-art on MESC and DFEW datasets while RL-driven trustworthiness improvements are validated through human and LLM evaluations, demonstrating its superior capability in applying a multimodal framework in this domain.
@@ 1 Introduction
Advances in large language models (LLMs) and vision–language models (VLMs) have transformed text generation, dialogue systems, and multimodal reasoning (NEURIPS2020_1457c0d6; mitsui2024pslm). These models now support applications in healthcare (aihealthcare; li2023llava; nguyen2024logra), summarization (summarizing), and retrieval (airetrieval; airetrieval2). Their versatility makes them promising tools for addressing complex social challenges, including mental health care (malgaroli2025large; obradovich2024opportunities). Systems such as Woebot show that AI-driven dialogue can alleviate depression and anxiety through cognitive behavioral therapy (CBT)–inspired conversations (woebot; textonly1; textonly2). However, most existing systems remain text-only, overlooking nonverbal cues—tone, facial expression, and gesture—that are essential for empathy and trust (multiemotion1). Empirical studies indicate that multimodal signals strengthen emotional understanding and engagement in human–computer interaction (multiemotion2; saffaryazdi2025empathetic; loveys2020effect; eyben2015geneva; wang2022systematic). Consequently, text-only systems often lack the authenticity and nuance required for effective emotional support.
@@ 2.3 Related works
Recent advancements in AI have enhanced emotional support systems, addressing limitations of smartphone-based conversational agents (smartphone; milne2022artificial), Muffin framework (muffin) uses model-agnostic AI feedback and contrastive learning to improve response fluency and relevance. Hybrid Empathetic Framework (HEF) (HEF) integrates LLMs with small-scale empathetic models to enhance emotion detection and response generation. The Sequential SMES framework (smes) leverages multimodal data to simulate therapeutic empathy and deliver tailored responses. Our MultiMood framework advances these efforts by incorporating trustworthiness through reinforcement learning with PPO and GRPO, ensuring safe, empathetic, and contextually appropriate responses for diverse user needs.
@@
Multimodal LLMs like InternVideo2.5 (internvideo2.5) and VideoLLaVA (videollava) advance video understanding. InternVideo2.5 employs a single InternVIT encoder with Hierarchical Token Compression (HICO) to efficiently process long videos, merging similar tokens to reduce computation while preserving quality (internvideo2.5). Its three-stage training supports tasks like temporal grounding and object tracking. VideoLLaVA aligns images and videos using LanguageBind for unified visual representation via a shared projection layer (videollava), but lacks audio processing, unlike MultiMood’s modality-specific projectors. These models focus on visual content while missing audio cues (volume, tone, pitch) critical for emotion recognition. Multimodal LLMs also enhance emotional support, overcoming single-modality limitations by capturing nuanced emotional signals for empathetic AI. EmotionLLaMA (emotionllama) uses the MERR dataset (28,618 samples) and specialized encoders for precise emotion recognition, surpassing GPT-4V in zero-shot tasks (zhang-etal-2024-visual). The SMES framework, with the MESC dataset (28,762 utterances from In Treatment), processes multimodal inputs for emotion recognition, strategy prediction, and response generation, improving therapeutic mimicry (smes). MultiMood stands out with specialized encoders per modality and a reinforcement learning algorithm designed for trustworthiness, as summarized in Table 1.
@@ A Spatial-Temporal Convolution (STC) connector (videollama2) captures spatial and temporal dynamics:
Our approach draws from ECoT (ECoT), a plug-and-play method that boosts LLM performance in emotional generation tasks by aligning with Goleman’s emotional intelligence theory (goleman1995emotional). We design a prompt template to guide the LLM in generating empathetic responses, integrating historical and real-time data. Multimodal features are concatenated into the input using specialized tokens: <video>, <audio>, and <history>, replaced by processed embeddings $E^{\prime}_{V}$, $E^{\prime}_{A}$, and $E^{\prime}_{H}$, respectively, forming the input sequence $X_{LLM}$. This attention-based fusion enables the model to dynamically prioritize relevant cues (e.g., tone, facial expressions) for safe, context-aware responses, while simultaneously predicting three classifications and generating therapist-like outputs (see Figure 2). More specific concatenation is discussed in the Appendix A.1.
@@ 4.1 Experimental Setup
We utilized the pretrained LLM from VideoLLaMA2 (videollama2) as the multimodal LLM backbone, leveraging its training on multimodal data. We compare MultiMood with API-based LLMs (GPT-4o (GPT4), Grok3 (xai2025grok2), Claude-3.7 (claude), Deepseek-R1 (deepseek), LLaMA4 (llama4)); Open-source LLMs (Qwen2 (qwen2) and Qwen2.5 (qwen25)) - finetuned on the MESC dataset (smes) with SFT and PPO; Multimodal frameworks (EmotionLLaMA (emotionllama), VideoLLaMA2, VideoLLaVA (videollava), InternVideo2.5-8B (internvideo2.5), VideoLLaMa3-7B (videollama3)) - not finetuned due to resource constraints; and SMES-leveraged models (MMGCN (MMGCN), MMDFN (MMDFN), Blenderbot (ESConv)).
@@ 7 Acknowledgments
This research was supported by The Mohamed bin Zayed University of Artificial Intelligence’s Travel Grant and the VNUHCM-University of Information Technology’s Scientific Research Support Fund.

### v2

# Reinforce Trustworthiness in Multimodal Emotional Support System

@@ Abstract
In today’s world, emotional support is increasingly essential, yet it remains challenging for both those seeking help and those offering it. Multimodal approaches to emotional support show great promise by integrating diverse data sources to provide empathetic, contextually relevant responses, fostering more effective interactions. However, current methods have notable limitations, often relying solely on text or converting other data types into text, or providing emotion recognition only, thus overlooking the full potential of multimodal inputs. Moreover, many studies prioritize response generation without accurately identifying critical emotional support elements or ensuring the reliability of outputs. To overcome these issues, we introduce MultiMood, a new framework that (i) leverages multimodal embeddings from video, audio, and text to predict emotional components and to produce responses responses aligned with professional therapeutic standards. To improve trustworthiness, we (ii) incorporate novel psychological criteria and apply Reinforcement Learning (RL) to optimize large language models (LLMs) for consistent adherence to these standards. We also (iii) analyze several advanced LLMs to assess their multimodal emotional support capabilities. Experimental results show that MultiMood achieves state-of-the-art on MESC and DFEW datasets while RL-driven trustworthiness improvements are validated through human and LLM evaluations, demonstrating its superior capability in applying a multimodal framework in this domain. The code for this paper is available at https://github.com/quangtuan-0504/Multimood.
@@ 1 Introduction
Advances in large language models (LLMs) and vision–language models (VLMs) have transformed text generation, dialogue systems, and multimodal reasoning (mitsui2024pslm). These models now support applications in healthcare (aihealthcare; nguyen2024logra), summarization (summarizing), and retrieval (airetrieval; airetrieval2). Their versatility makes them promising tools for addressing complex social challenges, including mental health care (malgaroli2025large). Systems such as Woebot show that AI-driven dialogue can alleviate depression and anxiety through cognitive behavioral therapy (CBT)–inspired conversations (woebot; textonly1). However, most existing systems remain text-only, overlooking nonverbal cues—tone, facial expression, and gesture—that are essential for empathy and trust (multiemotion1). Empirical studies indicate that multimodal signals strengthen emotional understanding and engagement in human–computer interaction (multiemotion2; saffaryazdi2025empathetic). Consequently, text-only systems often lack the authenticity and nuance required for effective emotional support.
@@ 2.3 Related works
Recent advancements in AI have enhanced emotional support systems, addressing limitations of smartphone-based conversational agents (smartphone), Muffin framework (muffin) uses model-agnostic AI feedback and contrastive learning to improve response fluency and relevance. Hybrid Empathetic Framework (HEF) (HEF) integrates LLMs with small-scale empathetic models to enhance emotion detection and response generation. The Sequential SMES framework (smes) leverages multimodal data to simulate therapeutic empathy and deliver tailored responses. Our MultiMood framework advances these efforts by incorporating trustworthiness through reinforcement learning with PPO and GRPO, ensuring safe, empathetic, and contextually appropriate responses for diverse user needs.
@@
Multimodal LLMs like InternVideo2.5 (internvideo2.5) and VideoLLaVA (videollava) advance video understanding. InternVideo2.5 employs a single InternVIT encoder with Hierarchical Token Compression (HICO) to efficiently process long videos, merging similar tokens to reduce computation while preserving quality (internvideo2.5). Its three-stage training supports tasks like temporal grounding and object tracking. VideoLLaVA aligns images and videos using LanguageBind for unified visual representation via a shared projection layer (videollava), but lacks audio processing, unlike MultiMood’s modality-specific projectors. These models focus on visual content while missing audio cues (volume, tone, pitch) critical for emotion recognition. Multimodal LLMs also enhance emotional support, overcoming single-modality limitations by capturing nuanced emotional signals for empathetic AI. EmotionLLaMA (emotionllama) uses the MERR dataset (28,618 samples) and specialized encoders for precise emotion recognition. The SMES framework, with the MESC dataset (28,762 utterances from In Treatment), processes multimodal inputs for emotion recognition, strategy prediction, and response generation, improving therapeutic mimicry (smes). MultiMood stands out with specialized encoders per modality and a reinforcement learning algorithm designed for trustworthiness, as summarized in Table 1.
@@ A Spatial-Temporal Convolution (STC) connector (videollama2) captures spatial and temporal dynamics:
Our approach draws from ECoT (ECoT), a plug-and-play method that boosts LLM performance in emotional generation tasks by aligning with Goleman’s emotional intelligence theory. We design a prompt template to guide the LLM in generating empathetic responses, integrating historical and real-time data. Multimodal features are concatenated into the input using specialized tokens: <video>, <audio>, and <history>, replaced by processed embeddings $E^{\prime}_{V}$, $E^{\prime}_{A}$, and $E^{\prime}_{H}$, respectively, forming the input sequence $X_{LLM}$. This attention-based fusion enables the model to dynamically prioritize relevant cues (e.g., tone, facial expressions) for safe, context-aware responses, while simultaneously predicting three classifications and generating therapist-like outputs (see Figure 2). More specific concatenation is discussed in the Appendix A.1.
@@ 4.1 Experimental Setup
We utilized the pretrained LLM from VideoLLaMA2 (videollama2) as the multimodal LLM backbone, leveraging its training on multimodal data. We compare MultiMood with API-based LLMs (GPT-4o (GPT4), Grok3 (xai2025grok2), Claude-3.7 (claude), Deepseek-R1 (deepseek), LLaMA4 (llama4)); Open-source VLMs (Qwen2 (qwen2) and Qwen2.5 (qwen25), EmotionLLaMA (emotionllama), VideoLLaMA2, VideoLLaVA (lin-etal-2024-video-llava), InternVideo2.5-8B (internvideo2.5), VideoLLaMa3-7B (videollama3)) - finetuned on the MESC dataset (smes) with SFT and PPO; VideoLLaMA2-72B and closed sources models are not finetuned due to resource constraints; and SMES-leveraged models (smes).
@@ 7 Acknowledgments
This research was supported by the Mohamed bin Zayed University of Artificial Intelligence Travel Grant. The authors also gratefully acknowledge the VNUHCM-University of Information Technology’s Scientific Research Support Fund for their financial assistance. The authors also thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Duy M. H. Nguyen. Duy M. H. Nguyen and Daniel Sonntag are also supported by the No-IDLE project (BMFTR, 16IW23002), the MASTER project (EU, 101093079), and the Endowed Chair of Applied Artificial Intelligence, Oldenburg University.